1. Margin of Safety

a. If Canace Company, with a break-even point at $256,000 of sales, has actual sales of $400,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number.

1. $

2.   %

b. If the margin of safety for Canace Company was 45%, fixed costs were $2,074,050, and variable costs were 55% of sales, what was the amount of actual sales (dollars)?
(Hint: Determine the break-even in sales dollars first.)
$

2.

Operating Leverage

Beck Inc. and Bryant Inc. have the following operating data:

Beck Inc. Bryant Inc.
Sales $286,600 $704,000
Variable costs (115,000) (422,400)
Contribution margin $171,600 $281,600
Fixed costs (105,600) (105,600)
Operating income $66,000 $176,000

a. Compute the operating leverage for Beck Inc. and Bryant Inc. If required, round to one decimal place.

Beck Inc.
Bryant Inc.

b. How much would operating income increase for each company if the sales of each increased by 20%? If required, round answers to nearest whole number.

Dollars Percentage
Beck Inc. $ %
Bryant Inc. $ %

3.

Break-Even Sales Under Present and Proposed Conditions

Portmann Company, operating at full capacity, sold 1,000,000 units at a price of $187 per unit during the current year. Its income statement is as follows:

Sales $187,000,000
Cost of goods sold (99,000,000)
Gross profit $88,000,000
Expenses:
Selling expenses $16,000,000
Administrative expenses 11,400,000
Total expenses (27,400,000)
Operating income $60,600,000

The division of costs between variable and fixed is as follows:

Variable Fixed
Cost of goods sold 70% 30%
Selling expenses 75% 25%
Administrative expenses 50% 50%

Management is considering a plant expansion program for the following year that will permit an increase of $11,220,000 in yearly sales. The expansion will increase fixed costs by $3,500,000 but will not affect the relationship between sales and variable costs.

Required:

1. Determine the total variable costs and the total fixed costs for the current year.

Total variable costs $
Total fixed costs $

2. Determine (a) the unit variable cost and (b) the unit contribution margin for the current year.

Unit variable cost $
Unit contribution margin $

3. Compute the break-even sales (units) for the current year.
units

4. Compute the break-even sales (units) under the proposed program for the following year.
units

5. Determine the amount of sales (units) that would be necessary under the proposed program to realize the $60,600,000 of operating income that was earned in the current year.
units

6. Determine the maximum operating income possible with the expanded plant.
$

7. If the proposal is accepted and sales remain at the current level, what will the operating income or loss be for the following year?
$
Answer #1

Margin of safety in dollars = Sales – Break even sales

= 400,000-256,000

= $144,000

% = Margin of Safety Sales/Total Sales

= 144,000/400,000

= 36%

b.Break even sales = Fixed costs/(1-Variable cost ratio)

= 2,074,050/(1-55%)

= $4,609,000

Actual Sales = Break even sales/(1- Margin of Safety ratio)

= 4,609,000/(1-45%)

= $8,380,000

a.Operating leverage = Contribution Margin/Operating income

Beck = 171600/66000 = 2.6

Bryant = 281600/176000

= 1.6

b.Increase in operating income = % increase in sales*Operating leverage

Beck = 20%*2.6 = 52% i.e. $34,320

Bryant = 20*1.6 = 32% i.e. $56,320